OpenStudy (anonymous):

Prove: ((cot^2-1)/csc^2) = cos2

OpenStudy (anonymous):

change into sin and cos

OpenStudy (anonymous):

see (cot^2 x - 1)/cosec^2x =((cos^2 x/ sin^2 x) - 1) / cosec^2x =( cos^2 x - sin^2 x)/sin^2 x ) / cosec^2x = ( cos^2 x - sin^2 x)/(sin^2 x * cosec^2x) =( cos^2 x - sin^2 x) =cos2x

OpenStudy (anonymous):

That was really, really, helpful! I didn't think of answer it in that manner. Thank you!

OpenStudy (anonymous):

welcome dear

OpenStudy (anonymous):

Last question, though! How did you get these: = ( cos^2 x - sin^2 x)/sin^2 x ) / cosec^2x = ( cos^2 x - sin^2 x)/(sin^2 x * cosec^2x) For the first one, is this what you did: ((cos^2/sin^2)*(1 or sin^2/sin^2))/csc^2? I'm trying to get it, but I can't figure out how you got the second one. :O

OpenStudy (anonymous):

( (cos^2 x - sin^2 x)/sin^2 x ) / (cosec^2x /1) = ( (cos^2 x - sin^2 x)/sin^2 x ) * (1/ cosec^2x) = ( cos^2 x - sin^2 x)/(sin^2 x * cosec^2x)